OpenStudy (anonymous):

A total of 32 students have entered a spelling contest. There are 6 medals for first through sixth place that will be awarded. In how many different ways can the medals be awarded?

OpenStudy (anonymous):

well @wio can u help this gye?

OpenStudy (anonymous):

Okay let me think

OpenStudy (anonymous):

You pick out on person from the group at a time.

OpenStudy (anonymous):

Since medals are distinct, you use permutation.

OpenStudy (anonymous):

In sort... \[ \large ^{32}P_6 = \frac{32!}{(32-6)!} = 32\times 31\times 30\times 29\times 28\times 27 \]